# AI Foundation Models in Remote Sensing: A Survey

### v1

@@ VI-A2 Practical Implications
  • Disaster Management. Models such as OFA-Net[84], DOFA [83] and Prithvi [33] are instrumental in flood mapping, as well as fire detection. These models provide critical real-time data that helps in identifying affected areas quickly, enabling timely and effective response measures. This capability supports emergency responds in prioritizing resource allocation and implementing evacuation plans, thereby reducing the impact of natural disasters. Additionally, these models assist in post-disaster recovery by assessing damage and monitoring the recovery process over time. By integrating various data sources, they enhance the ability to make informed decisions, coordinate response efforts, and plan for future disaster mitigation strategies.
VI-B Future Direction
@@
  • Efficient Model Development: Exploring techniques such as model distillation, pruning, and quantization to reduce computational requirements without compromising performance is crucial. Additionally, developing scalable architectures that efficiently handle ultra-high-resolution images is essential. For instance, applying pruning techniques to models like SatMAE [11] could maintain performance while reducing computational load.
@@
VI-C Limitation
@@
  • Scope and Coverage. The review focuses on foundation models released between June 2021 and June 2024. While the scope of this review is extensive and covers many significant developments, it is not exhaustive. Some recent advancements and innovations in the field may not be included due to their release timing or the lack of sufficient evaluation metrics at the time of writing. Consequently, certain cutting-edge models that have emerged in the latter part of this period or that have not yet been thoroughly evaluated might be omitted. This limitation underscores the need for readers to seek out the most current research and updates beyond the scope of this survey.
@@
VII Conclusion
In this comprehensive survey, we have reviewed the advancements in foundation models for remote sensing developed between June 2021 and June 2024. We categorized these models into vision and vision-language types, highlighting their unique methodologies and capabilities.
Our analysis covered various advanced techniques, including self-supervised learning (SSL), Vision Transformers (ViTs), and residual neural networks (ResNets). These models have significantly improved performance in tasks such as scene classification, semantic segmentation, and object detection, as well as in domain-specific applications like environmental monitoring, digital archaeology, agriculture, urban planning, and disaster management.
While significant progress has been made, several challenges persist, such as the need for more diverse and high-quality datasets, high computational requirements, and task-specific difficulties. Addressing these challenges will require further research and collaboration across disciplines.
@@ VII Conclusion
Acknowledgements. This work was supported by Vanderbilt Seeding Success Grant, Vanderbilt Discovery Grant, and VISR Seed Grant. We extend gratitude to NVIDIA for their support by means of the NVIDIA hardware grant. This works was also supported by NSF NAIRR Pilot Award NAIRR240055. This manuscript has been co-authored by ORNL, operated by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S.Department of Energy. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.

### v2

@@ VII-A2 Practical Implications
  • Disaster Management. Models such as OFA-Net[118], DOFA [117] and Prithvi [46] are instrumental in flood mapping, as well as fire detection. These models provide critical real-time data that helps in identifying affected areas quickly, enabling timely and effective response measures. This capability supports emergency responders in prioritizing resource allocation and implementing evacuation plans, thereby reducing the impact of natural disasters. Additionally, these models assist in post-disaster recovery by assessing damage and monitoring the recovery process over time. By integrating various data sources, they enhance the ability to make informed decisions, coordinate response efforts, and plan for future disaster mitigation strategies.
The improvements in accuracy across the models discussed have profound implications for real-world remote sensing applications. In deforestation monitoring, for instance, models like GFM achieve high pixel-level accuracy in semantic segmentation, showing up to a 4.5% improvement over baseline models, which enhances the precision of mapping forest cover changes, supporting conservation efforts[101]. Similarly, HyperSIGMA achieves an impressive 6.2% accuracy boost in hyperspectral vegetation monitoring, providing invaluable data for assessing forest health and biodiversity[95].
In urban planning, models like UPetu excel in infrastructure mapping by integrating multi-modal data, such as optical and radar imagery, achieving over 5% higher accuracy compared to single-modality models, which allows urban planners to make more informed land-use decisions[24]. Additionally, RingMo enhances object detection accuracy by 3.7% over traditional supervised models, effectively identifying dense urban features critical for disaster management and urban infrastructure assessment[87].
Finally, ORBIT demonstrates exceptional scalability, processing large climate datasets with a scaling efficiency of up to 85%, which supports applications in long-term environmental monitoring, such as climate change prediction and seasonal forecasting. This scalability not only advances traditional remote sensing workflows but also enables complex multi-temporal analyses and predictive modeling previously challenging with conventional methods[101].
While remote sensing has long benefited from multi-spectral and temporal data, the adaptability, scalability, and efficiency of foundation models unlock a new level of precision and accessibility in these applications. This advancement opens up opportunities to tackle complex and evolving challenges across domains—from environmental conservation to urban planning—that traditional models have struggled to address at scale.
VII-B Future Direction
@@
  • Efficient Model Development: Exploring techniques such as model distillation, pruning, and quantization to reduce computational requirements without compromising performance is crucial. Additionally, developing scalable architectures that efficiently handle ultra-high-resolution images is essential. For instance, applying pruning techniques to models like SatMAE [16] could maintain performance while reducing computational load. Model adaptation techniques such as LoRA (Low-Rank Adaptation) [41] have emerged as effective methods for fine-tuning large-scale models with minimal computational overhead. By decomposing weight updates into low-rank matrices, LoRA [41] enables efficient adaptation without the need to modify the entire set of model parameters, making it suitable for resource-constrained environments or when frequent re-training is required. Incorporating methods like LoRA [41] can further enhance the applicability of foundation models across diverse tasks and domains.
@@
Looking ahead, the consistent success of self-supervised learning methods in foundation models marks an exciting frontier for future research. These models’ ability to learn from unlabeled data and adapt to diverse remote sensing tasks with minimal fine-tuning suggests that advancements in unsupervised learning techniques could greatly reduce reliance on large labeled datasets, which remain a significant bottleneck in many remote sensing applications. However, as these models grow in size and complexity, balancing computational demands with the need for efficiency will become increasingly crucial. Future work may focus on developing more resource-efficient versions of foundation models that maintain high performance, particularly for deployment in real-time monitoring systems or environments with limited computational resources.
VII-C Limitation
@@
  • Scope and Coverage. The review focuses on foundation models released between June 2021 and June 2024. While the scope of this review is extensive and covers many significant developments, it is not exhaustive. Some recent advancements and innovations in the field may not be included due to their release timing or the lack of sufficient evaluation metrics at the time of writing. Consequently, certain cutting-edge models that have emerged in the latter part of this period or that have not yet been thoroughly evaluated might be omitted. This limitation underscores the need for readers to seek out the most current research and updates beyond the scope of this survey. Additionally, while foundation models have been empirically tested on a specific set of downstream applications, their robust architectures and general-purpose training paradigms, such as convolutional networks (e.g., ResNet) and vision transformers (e.g., ViT), indicate their potential to perform well across a much broader range of tasks. The limited testing observed in current literature should not be seen as a constraint on their applicability, but rather as an indication of the focus of existing research efforts. Given their design, these models are expected to generalize effectively to a wide variety of remote sensing tasks, even beyond those explicitly tested. Future work should aim to explore and validate their performance across more diverse applications to unlock their full potential.
@@
VIII Conclusion
In this comprehensive survey, we have reviewed the recent advancements in foundation models for remote sensing. We categorized these models based on their pretraining methods, image analysis techniques, and applications across different areas, highlighting their unique methodologies and capabilities.
Our analysis covered various advanced techniques, including self-supervised learning, vision transformers, and residual neural networks. These models have significantly improved performance on different image perception levels like region-level, pixel level, and image-level, as well as in applications like environmental monitoring, digital archaeology, agriculture, urban planning, and disaster management.
While significant progress has been made, several challenges persist, such as the need for more diverse and high-quality datasets, high computational requirements, and difficulties for different applications. Addressing these challenges will require further research and collaboration across disciplines.
@@ Acknowledgements
This work was supported by Vanderbilt Seeding Success Grant, Vanderbilt Discovery Grant, and VISR Seed Grant. We extend gratitude to NVIDIA for their support by means of the NVIDIA hardware grant. This works was also supported by NSF NAIRR Pilot Award NAIRR240055. This manuscript has been co-authored by ORNL, operated by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S.Department of Energy. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.
[Commonly Used Pre-train Dataset for Remote Sensing] RSD46-WHU [62, 116] dataset, introduced in 2017, is sourced from Google Earth and Tianditu. It contains 117,000 images with a patch size of 256 pixels and spatial resolutions ranging from 0.5 to 2 meters per pixel. Covering 46 categories globally, this dataset is primarily used for scene classification. Similarly, the Functional Map of the World (fMoW) [15], released in April 2018, comprises over 1 million images from Digital Globe. Spanning 63 categories across 207 countries, it includes multispectral images used for both scene classification and object detection.
SeasoNet [53], released in July 2022, features 1,759,830 images from Sentinel-2 with patch size from 20 to 120 pixels, supporting seasonal scene classification, segmentation, and retrieval over Germany. Lastly, the SSL4EO-S12 dataset [107] from November 2022 contains over 3 million images from Sentinel-1 and Sentinel-2, with patch size of 264x264 pixels. Since this dataset dose not contain any labels, it is commonly used for self-supervised learning.
In recent years, additional datasets have further enriched the resources available for remote sensing research. The SAMRS dataset [98], released in October 2023, offers a high-resolution collection of images sourced from datasets like HRSC2016 and FAIR1M-2.0, tailored for advanced segmentation tasks. With over 105,000 images and resolutions up to 1024x1024 pixels, SAMRS supports semantic and instance segmentation as well as object detection, contributing to the development of scalable segmentation models for remote sensing.
Focusing on change-aware learning, CACo [67], launched in June 2023, provides a variable patch-size dataset sourced from Sentinel-2. This dataset is optimized for change detection and contrastive learning, specifically addressing urban and rural landscapes. By prioritizing contrastive and self-supervised tasks, CACo aids in developing models that can adapt to changes in satellite imagery across various environments.
The SatlasPretrain [7] dataset, introduced in October 2023, is a large-scale collection with over 856,000 images combining Sentinel-2 and NAIP high-resolution sources. With multispectral and high-resolution imagery, SatlasPretrain supports applications such as land cover classification, segmentation, and change detection, further advancing research in high-resolution satellite image analysis.
The SSL4EO-L [83] dataset, released in October 2023, represents a vast resource with over 5 million images from Landsat, designed for self-supervised learning in cloud detection and land cover classification. By focusing on multi-year Landsat imagery, SSL4EO-L enables robust training for applications that benefit from long-term temporal coverage and cloud-resilient classification.
Finally, MMEarth [72], introduced in July 2024, combines data from Sentinel-1, Sentinel-2, and Aster DEM, providing over 1.2 million images for multimodal applications. This dataset supports land cover classification and semantic segmentation, enabling researchers to leverage multiple sensor types and climate data for improved geospatial representation learning.
! Month, Year Dataset Title Patch Size Size Resolution (m) Sensor Categories Geographic Coverage Image Type Application 2017 RSD46-WHU [62, 116] - 256 x 256 117,000 0.5 - 2 Google Earth, Tianditu 46 Global RGB Scene Classification Apr, 2018 fMoW [15] Functional Map of the World - 1,047,691 - Digital Globe 63 207 of 247 countries Multispectral Scene Classification, Object Detection May, 2019 DOTA [114] DOTA: A Large-scale Dataset for Object Detection in Aerial Images 800 × 800 to 20,000 × 20,000 11,268 Various Google Earth, GF-2 Satellite, and aerial images 18 Global RGB Object Detection Jun, 2019 SEN12MS [80] SEN12MS – A Curated Dataset of Georeferenced Multi-Spectral Sentinel-1/2 Imagery for Deep Learning and Data Fusion 256 x 256 541,986 10 Sentinel-1, Sentinel-2, MODIS Land Cover 33 Globally distributed SAR/Multispectral Land Cover Classification, Change Detection Jun, 2019 BigEarthNet [85] BigEarthNet: A Large-Scale Benchmark Archive For Remote Sensing Image Understanding 20 x 20 to 120 x 120 590,326 Various Sentinel-2 43 Europe Multispectral Scene Classification, Object Detection Jun, 2019 SeCo [66] Seasonal Contrast: Unsupervised Pre-Training from Uncurated Remote Sensing Data 264 x 264 $\sim$1M 10 - 60 Sentinel-2 - Global Multispectral Seasonal Change Detection, Land Cover Classification over Seasons Mar, 2021 MillionAID [64, 63] Million-AID 110 - 31,672 1,000,848 Various Google Earth 51 Global RGB Scene Classification Jul, 2021 Levir-KR [56] Geographical Knowledge-driven Representation Learning for Remote Sensing Images - 1,431,950 Various Gaofen1, Gaofen-2, Gaofen-6 8 Global Multispectral Change Detection, Scene Classification Apr, 2022 TOV-RS-Balanced [90] TOV: The Original Vision Model for Optical Remote Sensing Image Understanding via Self-supervised Learning 600 x 600 500,000 1 - 20 Google Earth 31 Global RGB Scene Classification, Object Detection, Semantic Segmentation Jul, 2022 SeasoNet [53] SeasoNet: A Seasonal Scene Classification, Segmentation and Retrieval dataset for satellite Imagery over Germany up to 120 x 120 1,759,830 10 - 60 Sentinel-2 33 Germany Multispectral Scene Classification, Scene Segmentation Nov, 2022 SSL4EO-S12 [107] SSL4EO-S12: A Large-Scale Multi-Modal, Multi-Temporal Dataset for Self-Supervised Learning in Earth Observation 264 x 264 3,012,948 10 - 60 Sentinel-1, Sentinel-2 - Global SAR/Multispectral Self-Supervised Learning Oct, 2023 SAMRS [98] SAMRS: Scaling-up Remote Sensing Segmentation Dataset with Segment Anything Model 600 x 600 to 1024 x 1024 105,090 Various HRSC2016, DOTA-V2.0, DIOR, FAIR1M-2.0 37 Global High-resolution Semantic Segmentation, Instance Segmentation, Object Detection Jun, 2023 CACo [67] Change-Aware Sampling and Contrastive Learning for Satellite Images Variable - 10 Sentinel-2 - Urban and Rural Areas Multispectral Semantic Segmentation, Change Detection, Self-Supervised Learning Oct, 2023 SatlasPretrain [7] SatlasPretrain: A Large-Scale Dataset for Remote Sensing Image Understanding 512 x 512 856,000 1 - 10 (Sentinel-2), 0.5 - 2 (NAIP) Sentinel-1, Sentinel-2, Landset, and NAIP 137 Global Multispectral, High-resolution Land Cover Classification, Segmentation, Change Detection Oct, 2023 SSL4EO-L [83] SSL4EO-L: Datasets and Foundation Models for Landsat Imagery 264 x 264 5,000,000 30 Landsat 4–5 TM, Landsat 7 ETM+, Landsat 8–9 OLI/TIRS - Global Multispectral Cloud Detection, Land Cover Classification, Semantic Segmentation Jul, 2024 MMEarth [72] MMEarth: Exploring Multi-Modal Pretext Tasks For Geospatial Representation Learning 128 x 128 1,200,000 10 Sentinel-2, Sentinel-1, Aster DEM 46 Global Multispectral, SAR, Climate Land Cover Classification, Semantic Segmentation
TABLE VIII: This table summarizes a set of commonly used pre-trained datasets for remote sensing, including details on dataset, sensor type, geographic coverage, and related applications.
Appendix A Biography Section
Siqi Lu (Student Member, IEEE) is a second-year Master’s student in Electrical and Computer Engineering at Vanderbilt University, supervised by Dr. Yuankai Huo and Dr. Mitchell M Wilkes. Her research interests include deep learning, medical image analysis, and software engineering. She received the B.S. degree in Electrical Engineering from the University of Illinois, Urbana-Champaign in 2023.
Junlin Guo is currently working toward the Ph.D. degree in Electrical and Computer Engineering at Vanderbilt University. His research interests include medical image analysis, deep learning, computer vision, and brain study. Before joining Vanderbilt University, he received the B.S. degree in telecommunication engineering from Northeastern University in 2017, and the M.S. degree from the Department of Electrical and Computer Engineering at Vanderbilt University in 2020, focusing on fMRI brain activation study.
James R Zimmer-Dauphinee received the B.A. degree in Anthropology and the B.S. degree in Mathematics from Georgia Southern University in 2011, the M.A. degree in Anthropology from the University of Arkansas in 2014, and the Ph.D. degree in Anthropology from Vanderbilt University in 2023. He is currently a postdoctoral fellow in the Spatial Analysis Research Laboratory (SARL) at Vanderbilt University, funded by the GeoPACHA 2.0 Grant from the National Endowment for the Humanities. His research interests include developing deep learning models for large-scale autonomous archaeological satellite imagery surveys, geophysical methods, and spatial modeling to understand the impact of colonization on indigenous peoples.
Jordan M Nieusma is a research assistant for the Vanderbilt University Spatial Analysis Research Laboratory. She received her M.S degree in Data Science at Vanderbilt University in May 2024 and holds a Bachelor of Arts degree in English with a French Minor from Haverford College.
Xiao Wang received the B.S. degrees in Mathematics and Computer Science from Saint John’s University, MN, in 2012, the M.S. degree in Electrical and Computer Engineering from Purdue University, West Lafayette, IN, in 2016, and the Ph.D. degree in Electrical and Computer Engineering from Purdue University in 2017. He pursued postdoctoral research at Harvard Medical School and Boston Children’s Hospital until 2021. He is currently a research staff scientist at Oak Ridge National Laboratory. His research interests include applying machine learning, medical physics, image processing, and high-performance computing to various imaging problems, including CT reconstruction, electron tomography imaging, and MRI. He was the 2022 AAPM Truth CT reconstruction challenge winner and a 2017 ACM Gordon Bell Prize finalist.
Parker VanValkenburgh is an Associate Professor of Anthropology and Interim Director of Latin American and Caribbean Studies at Brown University. His research focuses on the impacts of colonialism and imperialism on Indigenous people and environments in the Peruvian Andes. He utilizes diverse materials and digital methodologies, including GIS, to understand the transformation of relationships in imperial histories. He co-directs the Paisajes Arqueológicos de Chachapoyas (PACha) project and GeoPACHA (Geospatial Platform for Andean Culture, History, and Archaeology). VanValkenburgh received his Ph.D. from Harvard University.
Steven A Wernke is Associate Professor and Chair of Anthropology at Vanderbilt University, director of the Spatial Analysis Research Laboratory, and director of the Vanderbilt Institute for Spatial Research. Prof. Wernke is an archaeologist and historical anthropologist of the Andean region of South America. His research takes place at the intersection of several disciplines: archaeology and history, prehispanic and colonial studies, anthropology and cultural geography. Prof. Wernke’s interests center on the lived experiences of indigenous communities across the Spanish invasion of the Andes–especially how new kinds of communities, landscapes, and religious practice emerged out of successive attempts by the Inkas and the Spanish to subordinate and remake Andean societies in their own self-image. Methodologically, his work brings together analyses of archaeological and documentary datasets in geospatial frameworks.
Yuankai Huo is an Assistant Professor in Computer Science at Vanderbilt University, TN, USA. He received his B.S. degree in Electrical Engineering from Nanjing University of Posts and Telecommunications (NJUPT) in 2008, and Master degree in Electrical Engineering from Southeast University in 2011. After graduation, He worked in Columbia University and New York State Psychiatric Institute as a staff engineer and research officer from 2011 to 2014. He received his Master degree in Computer Science from Columbia University in 2014, and Ph.D. degree in Electrical Engineering from Vanderbilt University in 2018. Then, he had worked as a Research Assistant Professor at Vanderbilt University, and later, a Senior Research Scientist at PAII Labs. Since 2020, he has been a faculty member at the Department of Electrical Engineering and Computer Science, and Data Science Institute, Vanderbilt University.